OpenStudy (anonymous):

(-5a^3b^3c^0)^3 simplify

myininaya (myininaya):

so assuming c is not 0 then c^0=1

myininaya (myininaya):

\[(-5)^3(a^3)^3(b^3)^3=(-5)(-5)(-5)a^{3 \cdot 3}b^{3 \cdot 3} \] hope that helps :)